# Two-dimensional RMSD projections for reaction path visualization and validation

### 17 Feb 2026 v3

@@ 2 Results
We validate the framework, first contrasting against 1D profiles for a the 1,3-dipolar cycloaddition of ethylene and N2O forming 4,5-dihydro-1,2,3-oxadiazole, a well-studied benchmark reaction for NEB method development [20, 1, 3, 24, 29, 5]. Figure 1 compares conventional one-dimensional representations with the two-dimensional RMSD projection for energy-weighted NEB optimization in eOn [10] 222from https://eondocs.org using the PET-OMAT machine-learned potential [25, 4] using Metatomic [3]. The model has been trained on PBE reference data from the Open Materials dataset [2], with full construction details in the associated publication. For the cycloaddition, we contrast this with a saddle optimized from a different NEB calculation using the B3LYP functional in ORCA [1] projected onto the interpolated landscape.
The conventional profiles (Fig.1, bottom) show optimization from an initial barrier of approximately 1.1 eV to a final value near 0.4 eV over 120 steps, with the product lying approximately 0.8 eV below the reactant. The final path appears smooth and well-converged. These representations provide no information about sampling quality or the broader landscape topology.
The two-dimensional RMSD projection (Fig.1, top) provides some more insight. First, sampled structures (black dots) cluster tightly along the converged path, confirming robust single-pathway convergence. Second, the interpolated energy surface displays clear landscape topology: a barrier region (yellow/orange, approximately 0.4eV) separating the reactant basin (upper-left) from the deeper product basin (blue, lower-right). The projection reveals that the ORCA saddle configuration overlaps with the estimate from the MLIP.
@@
(b) Bicyclobutadiene Isomerization
@@
Figure 4: Two-dimensional RMSD projections of complex rearrangement pathways. (a) The Grignard rearrangement of (Z)-2-phenyl-2-((trimethylsilyl)oxy)hex-4-enenitrile exhibits a distinct ”turning” mechanism. The projection reveals significant optimization instability (scattered black dots) near the reactant basin that 1D profiles would obscure. The star indicates the reference saddle from Birkholz et al. (b) The conrotatory ring opening of bicyclobutadiene shows a sharp geometric ”kink” post-transition state. The projection confirms the machine-learned potential path (circles) faithfully traverses the saddle region identified by the Free String Method (star).
@@
To demonstrate robustness on non-trivial topologies, we analyze the Grignard rearrangement of (Z)-2-phenyl-2-((trimethylsilyl)oxy)hex-4-enenitrile [5] and the electrocyclic ring opening of bicyclobutadiene (Figure 2) [24]. Unlike the linear character of the cycloaddition, these reactions proceed through significant geometric turns. The Grignard pathway (Figure 2a) exhibits a curved trajectory where the path tangent rotates nearly 90 degrees in the $(RMSD_{R},RMSD_{P})$ plane. The 2D projection clearly captures the optimization noise (scattered black points) in the solvent-mediated reactant region, a diagnostic feature lost in scalar reaction coordinate plots, while demonstrating without explicitly needing to study the structures, the fact that they lie on the same energy contour. Similarly, the bicyclobutadiene isomerization (Figure 2b) displays a sharp geometric “elbow” immediately following the transition state. The interpolated surface correctly places the reference Free String Method (mlFSM) saddle [24] within the barrier region of the MLIP. This confirms that the NEB trajectory found the physically relevant saddle despite the complex topology, a validation that remains ambiguous when observing only the image index or scalar reaction coordinate 333or in practice, would involve normal mode analysis, barrier heights, RMSD, and visual inspection.
@@
To be explicit about what the 2D projection provides beyond standard scalar comparisons: computing energy differences and RMSD values between two saddle points tells one that they differ, but not how they differ relative to the reaction landscape. A 1D energy profile, by construction, collapses all geometric information onto the cumulative displacement axis, making it impossible to distinguish whether two saddle estimates from different methods lie on the same energy contour, sit in different basins, or whether the optimization explored alternative pathways before converging. The 2D projection preserves these geometric relationships. One can immediately see, for example, that the ORCA and MLIP saddles in Figure 1 occupy the same barrier region despite different path histories, or that the optimizer noise in the Grignard case (Figure 2a) is confined to a specific geometric region rather than distributed along the path. These are qualitative assessments, analogous to plotting a trajectory on a map rather than reporting only the total distance traveled.
@@
Furthermore, this framework provides a unified coordinate system for validating potential energy surfaces against higher-level theory. If these disagree, the relative differences between the saddles are apparent. The projections reveal whether resulting structures sit within comparable energy contours. Such an analysis, checking if the energy surface captures the qualitative barrier topology even if the precise saddle geometry differs due to functional sensitivity, is impossible with conventional one-dimensional profiles.
For complete reproduction, the materials archive includes both exact commands, and outputs at the time of submission.
@@
The $(r,p)$ projection is a lossy map from $\mathbb{R}^{3N}$ to $\mathbb{R}^{2}$. Multiple distinct Cartesian configurations can share the same RMSD coordinates, so the interpolated energy surface does not reconstruct the true PES but rather a projected slice of it. Apparent features far from the sampled data, for instance, alternative low-energy pathways visible in the color map, may not correspond to physical pathways in the full configuration space. The interpolation should be read as a qualitative guide near the data, not as a quantitative surrogate for the potential.
Because only the tangential force component enters the gradient projection, the method does not recover curvature information orthogonal to the path. In the full \$3N\$-dimensional PES, MEP images are by definition minima in the hyperplane orthogonal to the path. This property cannot be enforced or verified in the 2D projection without the Jacobian $\partial d_{\text{RMSD}}/\partial\mathbf{R}$, which is unavailable in closed form due to the rotation-permutation minimization. Any extension to orthogonal force projection would require differentiable structure matching, which is an open problem.
@@ Conclusion
In current practice, comparing NEB results across methods or potentials requires either pointwise structural inspection (impractical beyond a handful of images) or scalar summaries (barrier height, RMSD to reference) that discard geometric context. The 2D projection provides a middle ground: a single plot that preserves the geometric relationship between all sampled structures, the converged path, and any reference points, while remaining interpretable at a glance. The benchmarks demonstrate that this representation distinguishes numerical artifacts from physical topology and reveals whether different potentials produce qualitatively similar barrier regions, even when the precise saddle geometries differ.
While we focus on double-ended methods here, the $(RMSD_{R},RMSD_{P})$ projection operates independently of the path generation algorithm. One can project histories from single-ended saddle search methods, molecular dynamics, or meta-dynamics onto these intrinsic axes post hoc to diagnose path quality and basin hysteresis. The relative energy surfaces from multiple machine learning models could also be compared within such a framework.

### 24 Feb 2026 v4

@@ 2 Results
We empirically demonstrate the utility of the framework, first contrasting against 1D profiles for the 1,3-dipolar cycloaddition of ethylene and N2O forming 4,5-dihydro-1,2,3-oxadiazole, a well-studied benchmark reaction for NEB method development [21, 1, 3, 25, 30, 5]. Figure 3 compares conventional one-dimensional representations with the two-dimensional RMSD projection for energy-weighted NEB optimization in eOn [11] 222from https://eondocs.org using the PET-OMAT machine-learned potential [26, 4] using Metatomic [3]. The model has been trained on PBE reference data from the Open Materials dataset [2], with full construction details in the associated publication. For the cycloaddition, we contrast this with a saddle optimized from a different NEB calculation using the B3LYP functional in ORCA [1] projected onto the interpolated landscape.
The conventional profiles (Figure 3, bottom) show optimization from an initial barrier of approximately 1.1 eV to a final value near 0.4 eV over 120 steps, with the product lying approximately 0.8 eV below the reactant. The final path appears smooth and well-converged. These representations provide no information about sampling quality or the broader landscape topology.
The two-dimensional projection (Figure 3, top) provides additional insight. Sampled structures (black dots) cluster tightly along the converged path, confirming robust single-pathway convergence. The interpolated energy surface displays a barrier region (yellow/orange, approximately 0.4 eV) separating the reactant basin from the deeper product basin (blue). The projection reveals that the ORCA saddle configuration overlaps with the estimate from the MLIP. The posterior variance contours (dashed lines) confirm that the interpolation is tightly constrained near the path and visibly degrades in the periphery, serving as a built-in reliability indicator for the surface estimate.
@@
(b) Conrotatory Bicyclobutane Ring Opening
@@
Figure 4: Two-dimensional $(s,d)$ projections of complex rearrangement pathways with posterior variance contours (dashed lines). Both surfaces use the Nyström low-rank GP approximation; darker points denote the inducing subset (final converged NEB steps) while lighter points show the remaining optimization history retained for context. (a) The Grignard rearrangement of (Z)-2-phenyl-2-((trimethylsilyl)oxy)hex-4-enenitrile exhibits a distinct “turning” mechanism. The projection reveals optimization instability (scattered light dots) near the reactant basin that 1D profiles would obscure. The star indicates the reference saddle from Birkholz et al. (b) The conrotatory ring opening of bicyclobutane shows a sharp geometric “kink” post-transition state. The projection confirms the machine-learned potential path (circles) faithfully traverses the saddle region identified by the Free String Method (star).
@@
To assess robustness on non-trivial topologies, we analyze the Grignard rearrangement of (Z)-2-phenyl-2-((trimethylsilyl)oxy)hex-4-enenitrile [5] and the conrotatory ring opening of bicyclobutane (Figure 4) [25]. Unlike the linear character of the cycloaddition, these reactions proceed through significant geometric turns. The Grignard pathway (Figure 4(a)) exhibits a curved trajectory where the path tangent rotates nearly 90 degrees in the $(s,d)$ plane. The 2D projection captures the optimization noise (scattered lighter points) near the reactant basin, a diagnostic feature lost in scalar reaction coordinate plots, while showing without explicit structural inspection that the reference and computed saddles lie on the same energy contour. Similarly, the conrotatory bicyclobutane ring opening (Figure 4(b)) displays a sharp geometric “elbow” immediately following the transition state. The interpolated surface correctly places the reference Free String Method (mlFSM) saddle [25] within the barrier region of the MLIP. This confirms that the NEB trajectory found the physically relevant saddle despite the complex topology, a validation that remains ambiguous when observing only the image index or scalar reaction coordinate 333or in practice, would involve normal mode analysis, barrier heights, RMSD, and visual inspection.
@@
The 2D projection provides information beyond standard scalar comparisons. Computing energy differences and RMSD values between two saddle points tells one that they differ, but not how they differ relative to the reaction landscape. A 1D energy profile, by construction, collapses all geometric information onto the cumulative displacement axis. It cannot distinguish whether two saddle estimates from different methods lie on the same energy contour, sit in different basins, or whether the optimization explored alternative pathways before converging. The 2D projection preserves these geometric relationships. One can immediately see, for example, that the ORCA and MLIP saddles in Figure 3 occupy the same barrier region despite different path histories, or that the optimizer noise in the Grignard case (Figure 4(a)) is confined to a specific geometric region rather than distributed along the path. These are qualitative assessments, analogous to plotting a trajectory on a map rather than reporting only the total distance traveled.
@@
The framework also provides a unified coordinate system for validating potential energy surfaces against higher-level theory. If these disagree, the relative differences between the saddles are apparent. The projections reveal whether resulting structures sit within comparable energy contours. Determining whether the energy surface captures the qualitative barrier topology, even when the precise saddle geometry differs due to functional sensitivity, is impossible with conventional one-dimensional profiles.
For complete reproduction, the landscape figures are generated by a single command-line invocation. As an example, the Grignard projection (Figure 4(a)) is produced by:
python -m rgpycrumbs.cli --dev eon plt-neb \
--con-file neb.con --output-file jax_grad_imq.png \
--plot-type landscape --project-path \
--show-pts --landscape-path all \
--plot-structures crit_points \
--surface-type grad_imq_ny \
--ira-kmax 14 --show-legend \
--additional-con ts.xyz "Birkholz"
The --project-path flag (on by default) rotates the raw $(r,p)$ RMSD coordinates into the $(s,d)$ frame described above, --surface-type grad_imq_ny selects the Nystrom-accelerated gradient-enhanced IMQ kernel, --ira-kmax sets the adjacency cutoff distance for IRA graph-based structure matching, --plot-structures crit_points renders the reactant, product, saddle estimate, and highest-energy image, and --additional-con overlays a reference saddle from an external file. Cosmetic options (figure size, DPI, font size, rotation) are omitted for brevity. The materials archive includes exact commands and outputs at the time of submission.
@@
The $(r,p)$ projection is a lossy map from $\mathbb{R}^{3N}$ to $\mathbb{R}^{2}$. Multiple distinct Cartesian configurations can share the same RMSD coordinates, so the interpolated energy surface does not reconstruct the true PES but rather a projected slice of it. Apparent features far from the sampled data, for instance, alternative low-energy pathways visible in the color map, may not correspond to physical pathways in the full configuration space. The interpolation should be read as a qualitative guide near the data, not as a quantitative surrogate for the potential. The posterior variance contours overlaid on the figures directly visualize this degradation so that users can distinguish data-supported features from extrapolation artifacts.
As discussed in the methodology, only the tangential force component enters the gradient projection. The method does not recover curvature information orthogonal to the path, and the MEP orthogonality property cannot be enforced in the 2D projection. Any extension to orthogonal force projection would require differentiable structure matching. For molecular systems where the permutation group is large, this remains an open problem. For crystalline systems, where space group symmetry constrains the permutation set, tractable approximations may be feasible and are a natural direction for future work.
@@ Conclusion
In current practice, comparing NEB results across methods or potentials requires either pointwise structural inspection (impractical beyond a handful of images) or scalar summaries (barrier height, RMSD to reference) that discard geometric context. The 2D projection occupies the space between these extremes. A single plot preserves the geometric relationship between all sampled structures, the converged path, and any reference points, while remaining interpretable at a glance. The benchmarks show that this representation distinguishes numerical artifacts from physical topology and reveals whether different potentials produce qualitatively similar barrier regions, even when the precise saddle geometries differ. The rotated $(s,d)$ frame separates reaction progress from path deviation and gives a more physical coordinate system than raw RMSD distances. The posterior variance contours supply built-in reliability assessment for the interpolated surface. A Nystrom low-rank approximation extends the method to crystalline systems with large numbers of observations.
While we focus on double-ended methods here, the projection operates independently of the path generation algorithm. One can project histories from single-ended saddle search methods, molecular dynamics, or metadynamics onto these intrinsic axes post hoc to diagnose path quality and basin hysteresis. The relative energy surfaces from multiple machine learning models could also be compared within such a framework.
Acknowledgments
The author thanks Michele Ceriotti for suggesting the connection to the path collective variable formalism, and the members of Lab-COSMO at EPFL for discussions. The author thanks the anonymous reviewers for substantially improving the manuscript and the editor for facilitating the discussion. The author also thanks his family, pets, and plants for their patience and support. The author acknowledges the color scheme designed by Ruhila Goswami used for the figures.